OpenStudy (anonymous):

What is the perimeter of a rectangle with an area of 30 ft2 and a width of 5 ft (Answer:22ft? @Opcode ) 1 question to go.. and the other choices are 11 ft and 6 ft

OpenStudy (jack1):

heya @AngelicMelody so Area = 30 ft2 and Area = L x W so if width is 5 ft Area = L x W 30 = L x W 30 = L x 5 L = ... 30/5 =...?

OpenStudy (anonymous):

yeah I believe 22 is right

OpenStudy (jack1):

yep, perimeter is L + L + W + W = 6+6+5+5 = 22